Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (8,300 units) $ 282,200 $ 34.00
Variable expenses 157,700 19.00
Contribution margin 124,500 $ 15.00
Fixed expenses 55,000
Net operating income $ 69,500

Required:

(Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 50 units?

2. What would be the revised net operating income per month if the sales volume decreases by 50 units?

3. What would be the revised net operating income per month if the sales volume is 7,300 units?
Answer #1
Case 1 Case 2 Case 3
Units sold (8300+50)=8350 (8300-50)=8250 7300
Sales (8350*34)=283900 (8250*34)=280500 (7300*34)=248200
Variable expenses (8350*19)=158650 (8250*19)=156750 (7300*19)=138700
Contribution margin 125250 123750 109500
Fixed expenses 55,000 55,000 55,000
Revised net operating income $70250 $68750 $54500